# Gentle spaghettification in effective LQG dust collapse

### v1

@@ I Introduction
Spherically symmetric stellar collapse in classical general relativity generally leads to the formation of crush singularities, also called central singularities, according to the Penrose singularity theorem Penrose:1964wq . Even though such singularities have been widely studied in the literature and their mathematical properties analyzed in great detail, the common belief is that they are a pathological feature of Einstein theory of gravity, rather than a physical prediction. It has also been found that such singularities are not the only ones that develop during gravitational spherically symmetric collapse for generic initial density profiles. If neighboring layers of the collapsing star move at significantly different speeds, large inhomogeneity gradients can develop in the stellar energy density profile, sometimes leading to the so-called shell-crossing singularities (SCS) Penrose:1999vj ; Szekeres:1995gy ; Hellaby:1985zz ; Lasky:2006hq . Much less (though still significant) effort has been devoted to studying SCS compared to crush singularities, and the reasons are mainly threefold: firstly, during the collapse of pressureless matter, Hellaby and Lake proved Hellaby:1985zz that one can always choose initial data that do not develop such singularities before the central singularity is reached, which, in turn, is generally unavoidable. Additionally, by examining the mathematical features of these singularities, one can show Szekeres:1995gy that curvature scalars diverge less strongly compared to those associated with the crush singularities. Moreover, the behavior of matter falling toward these singularities is much less dramatic: despite the tidal forces being divergent at SCS, the angular and radial displacement of test particles remain finite Szekeres:1995gy . Furthermore, since the radial tidal force is negatively divergent, the so-called spaghettification effect is absent near SCS.
The purpose of this work is to understand how this picture changes once quantum gravitational corrections are taken into account. Among the possible candidate theories of quantum gravity, loop quantum gravity is one of the best-developed approaches Rovelli:2004tv ; Thiemann:2007pyv . Although a quantum gravitational model of stellar collapse is not yet available from the full theory, effective models have developed in recent years, where quantum gravitational corrections to Einstein equations are inspired by features of the underlying fundamental theory. The effective model on which this work is based is inspired from effective loop quantum cosmology, where the classical big bang singularity is replaced by a quantum gravitational bounce Ashtekar:2006wn . The modifications to Einstein field equations are based on the so-called $\bar{\mu}+K$-loop quantization scheme, where holonomy corrections with physical length $\sqrt{\Delta}$ (the minimum area gap in loop quantum gravity, proportional to the Planck length) affect only the angular component of the extrinsic curvature, and inverse triad corrections are neglected. The resulting effective dynamics has been extensively studied in recent years, both in the dust Husain:2021ojz ; Han:2023wxg ; Lewandowski:2022zce ; Giesel:2023hys ; Fazzini:2023scu ; Husain:2022gwp ; Kelly:2020lec ; Munch:2020czs ; Munch:2021oqn ; Cafaro:2024vrw ; Fazzini:2023ova ; Cipriani:2024nhx ; Bobula:2023kbo ; Bobula:2024chr ; Giesel:2023tsj and pressure cases Wilson-Ewing:2024uad ; Cafaro:2024lre . Despite much effort being devoted within this scheme to the Oppenheimer-Snyder collapse, the general picture regarding initial continuous decreasing inhomogeneous profiles is the following Fazzini:2023ova : the collapsing stellar core undergoes a quantum gravitational bounce when its energy density becomes planckian. Immediately after that, shell-crossing singularities arise, and the equations of motion break down. Although it is outside the scope of this work, for completeness, we mention that one can extend the dynamics beyond SCS formation by examining the integral form of the equation of motion, as in Painlevè-Gullstrand coordinates, it becomes a non-linear hyperbolic conservation law Leveque:2002 ; Batchelor_2000 , leading to the shockwave dynamics Husain:2022gwp ; Cipriani:2024nhx ; Fazzini:2025hsf . The same technique has been applied in the classical context Nolan:2003wp .
@@
The question this work aims to address is how tidal forces behave during this effective dynamics, focusing in particular on the bounce point and shell-crossing singularities. The main results we found are the following: angular tidal forces remain bounded during the bounce of the stellar core and reach their maximum exactly at the bounce point. Unlike the classical case, there is a short period of time around the bounce point in which the angular tidal force is positive, and two zeroes are attained, before and after the bounce. The radial component of the tidal force, instead, is also bounded at the bouncing point, but diverges at SCS. However, unlike the classical case, where the tidal forces negatively diverges at SCS, in the effective case, the radial component of the tidal force is positive at SCS, meaning that spaghettification of a test body occurs, similarly to what happens near the classical central singularity. Despite this, importantly, by solving the radial component of the geodesic deviation equation, one finds that the radial separation between two points remains finite at the shell-crossing singularity, meaning that the spaghettification does not result in an infinite stretching of the body. This allows us to conclude that such singularities, arising during effective stellar collapse, should not be considered as dramatic as the central classical singularities, and it motivates the extension of spacetime beyond these singularities through weak solutions.
@@ II Dust collapse in the marginally bound case
Effective dust collapse within the $\bar{\mu}+K-$loop quantization scheme has been extensively studied in Painlevè-Gullstrand coordinates Husain:2021ojz ; Kelly:2020lec ; Husain:2022gwp ; Fazzini:2023scu ; Cafaro:2024vrw ; Cipriani:2024nhx ; Fazzini:2025hsf and Lemaître-Tolman-Bondi coordinates (LTB) Giesel:2023hys ; Fazzini:2023ova ; Bobula:2024chr . The results derived in this work use LTB coordinates. The line element describing effective stellar collapse in these coordinates takes the following form
@@
where $\varepsilon$ is a constant of motion fixed by the initial condition, and determines the Newtonian energy (kinetic plus potential) of any shell $R$ at infinity. We will focus here on the marginally bound case $\varepsilon(R)=0$, $\forall R$. The function $r(R,t)$ is the areal radius of each shell $R$ of the distribution at time $t$, and its explicit form is provided by the solution of the effective equations. In turn, the equation of motion in LTB coordinates takes the following form Giesel:2023hys :
@@
where we defined $F(R)\equiv 2Gm(R)$ for compactness, and $m(R)$ is the (conserved) gravitational mass within the $R-$shell. As one can explicitly notice from (2), the spacetime evolution is provided by the evolution of the areal radius of the shells composing the collapsing star. The analytic solution of the previous equation reads:
@@
where $\alpha(R)$ is fixed by the initial condition. A fundamental feature of this generic solution, valid for any initial energy density profile, is that each shell $R$ will bounce at $t_{B}(R)=\alpha(R)$. In the pre-bounce (post-bounce) phase we have $t<\alpha(R)$ ($t>\alpha(R)$). Even though from (2) it seems that the shells’ dynamics is decoupled, and each shell evolves independently from the others (influenced only by the gravitational mass within its spherical volume), such solutions generally lead to shell-crossing singularities (SCS), at which points the equations cannot be trusted anymore. In order to define mathematically such singularities, we recall that the dust energy density in LTB coordinates reads
$\rho=\frac{\partial_{R}F}{8\pi Gr^{2}\partial_{R}r}~{}.$ (4)
A dynamical solution of the kind (3) develops SCS if, for a certain shell $R$, the following conditions are fulfilled: $\partial_{R}F\neq 0$, $\partial_{R}r(R,t)=0$. From (4), one can easily realize that these conditions imply a divergence in the dust energy density, and consequently a divergence in curvature scalars.
@@ III Choice of initial profiles
The effective equations (2) hold for any initial energy density dust configuration in the marginally bound case. In particular, one can easily derive the effective cosmological dynamics Ashtekar:2006wn by requiring $\rho(t_{0},R)=const.$, as well as the vacuum solution Giesel:2023hys ; Giesel:2024mps , by imposing $F=const$. Here, we are interested in stellar collapse, so we will assume that the initial energy density profile is positive and of compact support. Moreover, to keep the result as general as possible, we will consider decreasing inhomogeneous profiles ($\partial_{R}\rho(t_{0},R)<0$) of compact support, without specifying the initial data.
For these kind of profiles, we can combine a theorem, stated and proved in Fazzini:2023ova , and two results, one from Fazzini:2023ova and the other from Fazzini:2025hsf :
@@
    the latest time at which a shell-crossing singularity can arise in effective stellar collapse is $t=t_{B}(R)+(2/3)\sqrt{\Delta}$.
@@
    for initially decreasing continuous profiles, shell-crossing singularities cannot arise in a time $t\leq t_{B}(R)$.
Therefore, if we consider an initial profile that is non-negative, decreasing and of compact support, a shell-crossing singularity will necessarily form for a certain shell $R$ at time $t_{B}<t\leq t_{B}+(2\sqrt{\Delta}/3)$. These results allow us to restrict the analysis of tidal forces to the post-bounce dynamics of the shells, particularly to almost a planckian time interval after the bounce.
@@ IV Geodesic deviation equation for LTB collapse
In order to understand tidal effects during generic dust collapse, we need to study the geodesic deviation equation. In the classical theory, given $u^{\mu}$ the 4-vector tangent to the geodesics, and $\delta x^{\nu}$, the displacement vector orthogonal to $u$ ($\delta x^{\nu}u_{\nu}=0$), the Jacobi equation holds Wald:1984rg
@@
up to linear order in the deviation and its derivative. The covariant derivative is computed along the direction individuated by $u^{\mu}$, and therefore $\tau$ is the proper time along the geodesics. This equation does not uses the Einstein field equations but comes from the geometrical definition of the Riemann tensor. Therefore, it holds also at the effective level.
We are interested in the evolution of the deviation between dust particles during the stellar dynamics, and since dust particles follow geodesics also at the effective level (Appendix A), (5) describes tidal forces experienced by the dust itself. In particular, we are looking at a congruence of timelike geodesics for initially free-falling dust particles. Since the fluid $4$-velocity coincides with the $4$-velocity of a radially free-falling test particle, which in LTB coordinates takes the simple form $u^{\mu}=\{1,0,0,0\}$, equation (5) simplifies considerably
@@
In order to get rid of the covariant derivative and simplify the equation further, we use a common procedure and introduce tetrads dInverno:1992gxs . The tetrads associated with the LTB line element (1), which coincide with the tetrad basis in the free-falling frame, in the marginally bound case $\varepsilon(R)=0$ take the following (classical) form:
@@
These tetrads can be easily derived from $g^{\mu\nu}=e^{\mu}_{i}e^{\nu}_{j}\eta^{ij}$. Then, we write all the tensors appearing in (6) in terms of tetrads:
@@
where $a,b,c,d$ run from 00 to $3$. Notice that $\delta x^{0}=0$, since $u^{\mu}\delta x_{\mu}=0$ and $u^{\mu}$ has no spatial components. The deviation equation becomes
@@
which $0-$component is trivially satisfied. To arrive at (13) we used the tetrad postulate, which can be written as $De^{a}_{\mu}/D\tau=0$. The great advantage of using tetrads in LTB coordinates is that they define a locally inertial freely falling frame, for which the covariant derivative reduces to a simple ordinary derivative Szekeres:1995gy ; Fuchs(1990) . This leads to
@@
where the first equation determines the radial deviation and the index $i=2,3$ indicates the angular deviations in the tetradic basis. The time component of equation (15) gives a trivial identity. It is worth mentioning that, even though the functional dependence on $r$ takes the classical form, since the effective solution differs from the classical one, the same clearly holds for the solutions of equations (16) and (17).

### v2

@@ I Introduction
Spherically symmetric stellar collapse in classical general relativity generally leads to the formation of crush singularities, also called central singularities, according to the Penrose singularity theorem Penrose:1964wq . Even though such singularities have been widely studied in the literature and their mathematical properties analyzed in great detail, the common belief is that they are a pathological feature of Einstein theory of gravity, rather than a physical prediction. It has also been found that such singularities are not the only ones that develop during gravitational spherically symmetric collapse for generic initial density profiles. If neighboring layers of the collapsing star move at significantly different speeds, large inhomogeneity gradients can develop in the stellar energy density profile, sometimes leading to the so-called shell-crossing singularities (SCS) Penrose:1999vj ; Szekeres:1995gy ; Hellaby:1985zz ; Lasky:2006hq . Although shell-crossing singularities are also physical (i.e., not mere coordinate artifacts), much less (though still significant) effort has been devoted to studying SCS compared to crush singularities, and the reasons are mainly threefold: firstly, during the collapse of pressureless matter, Hellaby and Lake proved Hellaby:1985zz that one can always choose initial data that do not develop such singularities before the central singularity is reached, which, in turn, is generally unavoidable. Additionally, by examining the mathematical features of these singularities, one can show Szekeres:1995gy that curvature scalars diverge less strongly compared to those associated with the crush singularities. Moreover, the behavior of matter falling toward these singularities is much less dramatic: despite the tidal forces being divergent at SCS, the angular and radial displacement of test particles remain finite Szekeres:1995gy . Furthermore, since the radial tidal force is negatively divergent, the so-called spaghettification effect is absent near SCS.
The purpose of this work is to understand how this picture changes once quantum gravitational corrections are taken into account. Among the possible candidate theories of quantum gravity, loop quantum gravity is one of the best-developed approaches Rovelli:2004tv ; Thiemann:2007pyv . Although a quantum gravitational model of stellar collapse is not yet available from the full theory, effective models have developed in recent years, where quantum gravitational corrections to Einstein equations are inspired by features of the underlying fundamental theory. The effective model on which this work is based is inspired by effective loop quantum cosmology (LQC), where the classical big bang singularity is replaced by a quantum gravitational bounce Ashtekar:2006wn , arising when the energy density becomes planckian.
Modifications to Einstein’s equations at the effective level arise from a loop quantization of the midisuperspace, with spherical symmetry imposed at the classical level. In LQG-inspired models, such modifications typically affect only the geometrical part of the field equations, while the matter sector remains unaltered—a strategy that is also followed in the present work. This approach reflects a widely adopted methodology in Loop Quantum Cosmology (LQC), where the holonomy of the extrinsic curvature is quantized—referred to as $K$-loop quantization—in contrast to the full theory, where the holonomy of the entire Ashtekar-Barbero connection is quantized Bojowald:2003mc ; Vandersloot:2006ws ; Singh_2013 . In this treatment, inverse triad corrections are neglected. The physical length over which the holonomy is computed is determined via the improved dynamics scheme (the so-called $\bar{\mu}$-scheme), which ensures that quantum corrections are suppressed in the low-curvature regime Ashtekar:2006wn and connects the effective theory to the minimum non-zero area eigenvalue of the full quantum theory $\Delta\propto l_{Pl}^{2}$.
This quantization prescription has been implemented explicitly in Husain:2022gwp , where both the areal gauge and dust-time gauge are fixed at the classical level. However, the dynamics is analyzed only at the effective level. The expectation that these effective equations reliably capture the underlying quantum dynamics stems from an argument inherited from LQC. Specifically, it has been shown Rovelli:2013zaa that for cosmological solutions with physical volumes at the bounce satisfying $V_{phys}\gg\Delta^{\frac{3}{2}}$ , quantum fluctuations in the gravitational field become negligible even when the curvature approaches the Planck scale. This explains why quantum and effective cosmological dynamics agree so well numerically when large physical volumes are considered.
To apply the same reasoning to gravitational collapse, consider the simplest case: Oppenheimer–Snyder (OS) collapse. The interior solution of the OS model is isomorphic to a cosmological spacetime, and the physical volume of the star at the bounce is much larger than Planckian Husain:2021ojz . Therefore, quantum fluctuations are also expected to be negligible in this astrophysical context. It is important to note, however, that this argument may not hold in the case of inhomogeneous collapse, where large inhomogeneities in the energy density develop—the scenario explored in this work. A comment on this limitation is provided in the concluding discussion. Motivated by feasibility and also by the previous argument, the effective dynamics has been extensively studied in recent years, both in the dust Husain:2021ojz ; Han:2023wxg ; Lewandowski:2022zce ; Giesel:2023hys ; Fazzini:2023scu ; Husain:2022gwp ; Kelly:2020lec ; Munch:2020czs ; Munch:2021oqn ; Cafaro:2024vrw ; Fazzini:2023ova ; Cipriani:2024nhx ; Bobula:2023kbo ; Bobula:2024chr ; Giesel:2023tsj and pressure cases Wilson-Ewing:2024uad ; Cafaro:2024lre . For a comparison with other existing effective models, see Ashtekar:2023cod .
Despite much effort being devoted within this scheme to the Oppenheimer-Snyder collapse, where the stellar dynamics follows exactly the cosmological one Munch:2020czs ; Munch:2021oqn ; Lewandowski:2022zce ; Fazzini:2023scu ; Giesel:2023hys the general picture regarding initial continuous decreasing inhomogeneous profiles is different Fazzini:2023ova : the collapsing stellar core undergoes a quantum gravitational bounce when its energy density becomes planckian. Immediately after that, shell-crossing singularities arise (which, as mentioned before, are physical singularities), and the equations of motion break down. Shell-crossing singularities produce a non-isolated thin shell of matter, which dynamics can in principle be studied with a dynamical version of the Israel junction conditions Fazzini:2025zrq , a tool that is not yet available.
@@
The question this work aims to address is how tidal forces behave during this effective dynamics, focusing in particular on the bounce point and shell-crossing singularities. The main results we found are the following: the angular tidal forces remain bounded during the bounce of the stellar core and reach their maximum exactly at the bounce point. Unlike the classical case, there is a short period of time around the bounce point, in which the angular tidal force is positive, and two zeroes are attained, just before and after the bounce. The radial component of the tidal force, instead, is also bounded at the bouncing point but diverges at SCS. However, unlike the classical case, where the tidal forces negatively diverge at SCS, in the effective case, the radial component of the tidal force is positive at SCS, meaning that spaghettification of a test body occurs, similarly to what happens near the classical central singularity. Despite this, importantly, by solving the radial component of the geodesic deviation equation, one finds that the radial separation between two points remains finite at the shell-crossing singularity, meaning that the spaghettification does not result in an infinite stretching of the body. This allows us to conclude that such singularities, arising during effective stellar collapse, should not be considered as dramatic as the central classical singularities, and it motivates the extension of spacetime beyond these physical singularities.
@@ II Dust collapse in the marginally bound case
Effective dust collapse within the $\bar{\mu}+K-$loop quantization scheme has been extensively studied in Painlevè-Gullstrand coordinates Husain:2021ojz ; Kelly:2020lec ; Husain:2022gwp ; Fazzini:2023scu ; Cafaro:2024vrw ; Cipriani:2024nhx ; Fazzini:2025hsf and Lemaître-Tolman-Bondi coordinates (LTB) Giesel:2023hys ; Fazzini:2023ova ; Bobula:2024chr .
Although the original formulation of the model—both at the quantum and effective levels—was carried out within the dust-time and areal gauges, it has been shown that the areal gauge condition can be relaxed at the classical level, and imposed directly at the effective level. This leads to an effective model featuring an unmodified classical diffeomorphism constraint, in line with the structure of the full theory Giesel:2023hys . This relaxation enables the study of dynamics in various spatial gauges at the effective level, with the effective equations of Husain:2022gwp recovered upon imposing the areal gauge. The flexibility in choosing the spatial gauge at the effective level is particularly valuable when analyzing the formation and physical properties of shell-crossing singularities Fazzini:2023ova . For this reason, the results presented in this work make use of LTB coordinates.
The line element describing effective stellar collapse in these coordinates takes the following form
@@
where $\varepsilon$ is a time-independent function fixed by the initial condition, and determines the Newtonian energy (kinetic plus potential) of any shell $R$ at infinity. We will focus here on the marginally bound case $\varepsilon(R)=0$, $\forall R$. The function $r(R,t)$ is the areal radius of each shell $R$ of the distribution at time $t$, and its explicit form is provided by the solution of the effective equations written in the LTB gauge. In turn, the equation of motion in LTB coordinates takes the following form Giesel:2023hys :
@@
where we defined $F(R)\equiv 2Gm(R)$ for compactness, and $m(R)$ is the ADM mass function for the portion of the star within the $R$-shell. The factor $\gamma$ is the Barbero-Immirzi parameter, that for simplicity we set to 111. As one can explicitly notice from (2), the spacetime evolution is provided by the evolution of the areal radius of the shells composing the collapsing star. The analytic solution of the previous equation reads:
@@
where $\alpha(R)$ is fixed by the initial condition. A fundamental feature of this generic solution, valid for any initial energy density profile, is that each shell $R$ will bounce at $t_{B}(R)=\alpha(R)$. In the pre-bounce (post-bounce) phase, we have $t<\alpha(R)$ ($t>\alpha(R)$). Even though from (2) it seems that the shells’ dynamics is decoupled, and each shell evolves independently from the others (influenced only by the gravitational mass within its spherical volume), such solutions generally lead to shell-crossing singularities, at which point the equations cannot be trusted anymore. In order to define mathematically such singularities, we recall that the dust energy density in LTB coordinates reads Giesel:2023hys
$\rho=\frac{\mathcal{H}_{dust}}{4\pi\sqrt{q}}=\frac{\partial_{R}F}{8\pi Gr^{2}\partial_{R}r}~.$ (4)
Here, $\sqrt{q}$ denotes the determinant of the spatial metric (1), and $\mathcal{H}_{dust}$ is the dust Hamiltonian density, integrated over ${\rm d}\theta$ and ${\rm d}\phi$, while $\rho$ represents the dust energy density. Note that these expressions retain their classical functional form, since the matter sector is kept classical in this effective model Kelly:2020lec . As in the classical case, the LTB coordinates are adapted to the comoving dust frame. Therefore, $\rho$ corresponds to the $tt$ component of the energy-momentum tensor. Note that the functional relation between the lagrangian density for dust and the energy-momentum tensor is classical. Although the functional form of (4) remains classical, its explicit behavior deviates from the classical solution when $\rho$ approaches the critical density $\rho_{crit.}=3/(8\pi G\Delta)\sim\rho_{\text{Pl.}}$, or equivalently when $r(R,t)\sim[F(R)\Delta]^{1/3}$—that is, near the bounce.
A dynamical solution of the form (3) develops a SCS if, for a given shell labeled by $R$, the following conditions are satisfied: $\partial_{R}F\neq 0$ and $\partial_{R}r(R,t)=0$. From equation (4), it is easy to see that these conditions lead to a divergence in the dust energy density, which in turn results in a divergence of the curvature scalars Giesel:2024mps .
@@ III Choice of initial profiles
The effective equations (2) hold for any initial energy density dust configuration in the marginally bound case. In particular, one can easily derive the effective cosmological dynamics Ashtekar:2006wn by requiring $\rho(t_{0},R)=const.$, as well as the vacuum solution Giesel:2023hys ; Giesel:2024mps , by imposing $F=const$. Here, we are interested in stellar collapse, so we will assume that the initial energy density profile is positive and has compact support. Moreover, to keep the result as general as possible, we will consider decreasing inhomogeneous profiles ($\partial_{R}\rho(t_{0},R)<0$) of compact support, without specifying the initial data.
For these kinds of profile, we can combine a theorem, stated and proved in Fazzini:2023ova , and two results, one from Fazzini:2023ova and the other from Fazzini:2025hsf :
@@
    the latest time at which a shell-crossing singularity can arise in effective stellar collapse is $t=t_{B}(R)+(2/3)\sqrt{\Delta}$ Fazzini:2023ova .
@@
    for initially decreasing continuous profiles, shell-crossing singularities cannot arise in a time $t\leq t_{B}(R)$ Fazzini:2025hsf .
Therefore, if we consider an initial profile that is non-negative, decreasing, and of compact support, a shell-crossing singularity will necessarily form for a certain shell $R$ at time $t_{B}<t\leq t_{B}+(2\sqrt{\Delta}/3)$. These results allow us to restrict the analysis of tidal forces to the post-bounce dynamics of the shells, particularly to almost a planckian time interval after the bounce.
@@ IV Geodesic deviation equation for LTB collapse
In order to understand the tidal effects during generic dust collapse, we need to study the geodesic deviation equation. In the classical theory, given $u^{\mu}$ the 4-vector tangent to the geodesics, and $\delta x^{\nu}$, the displacement vector orthogonal to $u$ ($\delta x^{\nu}u_{\nu}=0$), the Jacobi equation holds Wald:1984rg
@@
up to linear order in the deviation and its derivative. The covariant derivative is computed along the direction individuated by $u^{\mu}$, and therefore $\tau$ is the proper time along the geodesics. This equation does not use the field equations (classical or modified), but comes from the geometrical definition of the Riemann tensor. Therefore, it also holds at the effective level.
We are interested in the evolution of the deviation between dust particles during the stellar dynamics, and since dust particles follow geodesics also at the effective level (which can be readily verified by checking that the dust four-velocity $u^{\mu}=(1,0,0,0)$ satisfies the geodesic equation), (5) describes tidal forces experienced by the dust itself. In particular, we are looking at a congruence of time-like geodesics for initially free-falling dust particles. Since the fluid 444-velocity coincides with the 444-velocity of a radially free-falling test particle, equation (5) simplifies considerably
@@
In order to get rid of the covariant derivative and further simplify the equation, we use a common procedure and introduce tetrads dInverno:1992gxs . The tetrads associated with the LTB line element (1), which coincide with the tetrad basis in the free-falling frame, in the marginally bound case $\varepsilon(R)=0$ take the following (classical) form:
@@
These tetrads can easily be derived from $g^{\mu\nu}=e^{\mu}_{i}e^{\nu}_{j}\eta^{ij}$. Then, we write all the tensors appearing in (6) in terms of tetrads:
@@
where $a,b,c,d$ run from 0 to 333. Notice that $\delta x^{0}=0$, since $u^{\mu}\delta x_{\mu}=0$ and $u^{\mu}$ have no spatial components. The deviation equation becomes
@@
which 00-0 -component is trivially satisfied. To arrive at (13) we used the tetrad postulate, which can be written as $De^{a}_{\mu}/D\tau=0$. The great advantage of using tetrads in LTB coordinates is that they define a locally inertial freely falling frame, for which the covariant derivative reduces to a simple ordinary derivative Szekeres:1995gy ; Fuchs(1990) . This leads to
@@
where the first equation determines the radial deviation and the index $i=2,3$ indicates the angular deviations in the tetradic basis. The time component of (15) gives a trivial identity. It is worth mentioning that even though the functional dependence on $r$ takes the classical form, since the effective solution differs from the classical one, the same clearly holds for the solutions of equations (16) and (17).